OpenStudy (anonymous):

A golfer imparts a speed of 30.3 m/s to a ball, and it travels the maximum possible distance before landing on the green. The tee and the green are at the same elevation. (a) How much time does the ball spend in the air? (b) What is the longest hole in one that the golfer can make, if the ball does not roll when it hits the green?

OpenStudy (shamim):

Range will b maximum when the angle of projection is 45°

OpenStudy (shamim):

At maximum height velocity v=0 Right?

OpenStudy (shamim):

Now u hv to use the formula v=u-gt 0=30.3*sin45-9.8t t=?

OpenStudy (shamim):

T=2*t

OpenStudy (shamim):

Feel free to ask for more clarification